Circular No. 7505

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000dr IN IC 1610 M. Modjaz, D. Weisz, and W. D. Li, University of California at Berke-

ley, on behalf of the Lick Observatory Supernova Search (cf. IAUC 6627, 7126), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova on unfiltered images taken on Oct. 5.3 (mag ∼ 18.1) and 6.3 UT (mag ∼ 17.8). The new object is located at α = 1^{h}01^{m}44^{s}.04, δ = −15^{o}34′09′′.1 (equinox 2000.0), which is 20′′.9 east and 5′′.5 south of the nucleus of IC 1610. A KAIT image of the same field taken on Sept. 27.3 UT showed nothing at the position of SN 2000dr (limiting mag ∼ 19.0).

SUPERNOVA 2000dp IN NGC 1139 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report: “A spectrum of SN 2000dp (cf. IAUC 7503), obtained by P. Berlind on Oct 4.38 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-Ia supernova at maximum light. The spectrum exhibits a blue continuum, with well-defined Si II, S II, Fe II, Mg II, and Ca II features. From a separate exposure of the nucleus, we derive a recession velocity for NGC 1139 of 10 350 km/s. The supernova expansion velocity is 12 700 km/s for Si II (rest 635.5 nm). The spectral feature age of the supernova (Riess et al. 1997, A.J. 114, 722) is 0 ± 2 days before maximum light.”

A. V. Filippenko and R. Chornock, University of California at Berkeley, report that inspection of CCD spectra (range 330–780 nm) of SN 2000dp obtained on Oct. 6 UT with the Shane 3-m reflector at Lick Observatory reveals that the object is a normal type-Ia supernova, 3 ± 2 days past maximum brightness (using the spectral feature aging method of Riess et al., op. cit.). Cross-correlation of the spectrum with that of type-Ia SN 1994D at a similar epoch yields a redshift (cz) of ∼ 10 000 km/s.

COMET C/1999 T1 (McNAUGHT-HARTLEY) Total visual magnitude estimates: Sept. 23.85 UT, 9.9 (A. Pearce,

Nedlands, W. Australia, 0.20-m reflector); Oct. 2.71, 9.6 (D. A. J. Seargent, The Entrance N.S.W., 25×100 binoculars); 4.81, 9.8 (Pearce); 6.29, 10.0: (A. Amorim, Florianopolis, Brazil, 0.14-m reflector).

2000 October 7  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green